# MindChat: A Privacy-preserving Large Language Model for Mental Health Support

### v1

@@ 4 Experiments
4.1 Baselines
For the dataset-level comparison, we evaluate MindCorpus against seven existing emotional dialogue datasets. These include four Chinese datasets: PsyDTCorpus xie2025psydt, SoulChatCorpus chen-etal-2023-soulchat, SMILECHAT qiu-etal-2024-smile, and CPsyCounD zhang-etal-2024-cpsycoun, as well as three English datasets: ExTES zheng-etal-2024-self, ESD-CoT zhang-etal-2024-escot, and AUGESC zheng-etal-2023-augesc.
At the model level, MindChat is benchmarked against a diverse set of LLMs. This set includes widely used general-purpose models such as ChatGPT achiam2023gpt, DeepSeek liu-2024-deepseek, and Gemini team-2024-gemini, the base model Qwen3-8B yang2025qwen3, and four specialized mental health LLMs SoulChat2.0 chen-etal-2023-soulchat, CPsyCounX zhang-etal-2024-cpsycoun, EmoLLM2.0 yang2024emollm, and MeChat qiu-etal-2024-smile.
4.2 Experimental settings
@@
Figure 4: Data distribution of MindCorpus.
Table 2: Basic statistics of emotional dialogue datasets.
@@
4.3 Results
4.3.1 Synthesized data analysis
The synthesized dataset MindCorpus contains 5.75.7k dialogue sessions across diverse themes, generated from curated scenario scripts using a multi-agent collaborative framework. As shown in Fig. 4, dialogues are distributed across multiple themes. On average, each session comprises $12.0$ utterances, with each utterance having an average length of $84$ words. Table 2 summarizes the basic statistics of MindCorpus in comparison with existing emotional dialogue datasets.
Table 3: Automatic and human evaluations of emotional dialogue datasets. Highest scores are in bold.
@@ 4.3.1 Synthesized data analysis
Table 4: Spearman rank correlations between automatic and human evaluations across emotional dialogue datasets, reported per evaluation dimension.
@@
To evaluate data quality, we follow the sampling protocol of zheng-etal-2024-self, randomly selecting $50$ sessions from each dataset while ensuring diversity by incorporating data from clearly classified content categories. We employ GPT-4o as a judge model to score dialogues based on our five-dimensional evaluation framework. Table 3 reports the results of both automated and human evaluations. The results of automatic evaluation show that MindCorpus attain the best performance on all metrics except for Tru. In human evaluation, aside from Emo., the remaining four metrics consistently indicate that MindCorpus outperforms other emotional dialogue datasets. Overall, both automated and human evaluations confirm the superior quality of MindCorpus compared to existing emotional dialogue datasets. To further examine the consistency between automatic and human evaluations, Spearman rank correlations zar2005spearman are computed for each evaluation dimension at the dataset level. As shown in Table 4, most dimensions exhibit moderate to high agreement, demonstrating strong consistency between automated and human evaluations and supporting the reliability of the reported results in Table 3.
4.3.2 Heterogeneous multi-LLM coordination
We examine the performance of heterogeneous LLMs collaborating within the proposed multi-agent framework, including models of varying sizes and different developers. Table 5 presents the performance of these LLMs working together in $11$ coordination groups.
Table 5: Evaluation results for data synthesized by heterogeneous multi-agent groups. Highest scores are in bold, second-highest are underlined.
@@ 4.3.2 Heterogeneous multi-LLM coordination
During the experiments, we observe that excessively long model outputs tend to exhibit content homogenization, which deviates from real-world psychological counseling scenarios. Empirical statistics of publicly available emotional and counseling dialogue datasets in Table 2 indicate that the average length of a single response is generally around or below $100$ words. Motivated by this distribution, we treat dialogues with single-round response lengths under $100$ words as qualified data. Accordingly, Groups $6$ to $11$ in Table 5 meet this criterion.
With different collaboration groups, the proportion of dialogues requiring modification changes, implying various interactions between agents. More interactions typically result in longer durations. As shown in Groups 5 and 6, compared to Qwen2.5-3B-Instruct, MiniCPM-4B produces fewer qualified dialogues and generates relatively longer responses. However, GPT-4o-based evaluation scores are higher for longer responses, possibly because the automated judges evaluate responses using point-based criteria, whereby longer outputs are more likely to satisfy multiple scoring aspects. In general, larger model sizes correlate with higher data quality, though this trend is not strictly monotonic. In Groups 10 and 11, we employ commercial models for data synthesis via API calls. The synthesized data quality in Group 11 surpasses that of all other experimental groups. However, Group 10 underperforms relative to the small-model collaborations in Groups 8 and 9. This discrepancy may be attributed to performance differences between DeepSeek-V3 and Qwen-Max. As shown in Table 5, the quality of data generated through multi-agent collaboration depends not only on model scale but also strongly on model type. Notably, relatively smaller-scale models also demonstrate the potential to achieve performance comparable to that of commercial-grade models.
@@
4.3.3 Impact of key agents on data quality
@@
4.3.4 Comparison of different psychological LLMs
@@
In order to evaluate the performance differences between MindChat and other benchmark models, we conduct both automatic and human evaluations. The results are summarized in Table 6. In the automatic evaluation, MindChat achieves the highest scores across all metrics, better than the general LLMs and the existing mental health-focused LLMs. It is worth noting that MindChat demonstrates superior performance in the dimensions of Comprehensiveness, Professionalism, and Authenticity. These three metrics reflect the depth and quality of psychological support provided, thus highlighting the enhanced capability of MindChat in addressing the psychological background of seeker, offering professional guidance, and establishing authentic empathetic connections. In the human evaluation, DeepSeek and Gemini, as general LLMs with significantly larger parameter sizes, occupy the top two respectively. Despite this disadvantage in terms of model scale, MindChat still ranks third overall. More importantly, it significantly outperforms other models of comparable size, including both general-purpose and domain-specific psychological LLMs, underscoring its effectiveness and competitiveness in mental health scenarios while preserving user privacy.
To further assess the consistency between automatic and human evaluations at the model level, Spearman rank correlations are computed for each evaluation dimension. The results show that Comprehensiveness achieves a high correlation of $0.730$ with a p-value of $0.026$, while Professionalism and Authenticity exhibit lower correlations of $0.351$ with a p-value of $0.354$ and $0.378$ with a p-value of $0.316$, respectively. The observed correlations show clear consistency between automated and human evaluations in Comprehensiveness, whereas Professionalism and Authenticity exhibit relatively weaker alignment. By offering complementary yet distinct perspectives, the two evaluation schemes jointly strengthen the credibility of the model comparison results reported in Table 6.
Figure 7: Comparison of slices: (a) Same theme, (b) Different themes.
Table 6: Automatic and human evaluations of multiple LLMs in psychological counseling capabilities.
@@
Table 7: Impact of data quantity and diversity on the performance of MindChat.
@@
4.3.5 Effects of data quantity and diversity
In FL architecture, an increasing number of participating clients often leads to larger data quantities and potentially greater diversity. This section investigates how the scale and diversity of training data affect the performance of MindChat under the designed training framework. To evaluate the impact of data quantity, we sample 600 sessions from the Depression and Anxiety dataset and partition them into six equal subsets, as shown in LABEL:subfig:quantity. For data diversity analysis, we collect 100 samples from each of six distinct themes, outlined in LABEL:subfig:diversity. All experiments are conducted with $\epsilon=1$, and model performance is compared across settings with $2$, 44, and $6$ collaborating clients. A local baseline is trained using $100$ samples from a single theme.
As shown in Table 7, FL enhances model generalization ability under data isolation, demonstrating the benefits of multi-party collaboration. When training within the same theme, performance increases with the number of clients, indicating the benefit of larger data volume. A similar trend is observed in cross-theme settings, where both data quantity and thematic diversity contribute to performance gains. Notably, models trained with diverse themes outperform those trained with increased data quantity from a single theme. For instance, MindChat-T2, which uses two distinct themes, achieves a higher average score than MindChat-Q6, trained on six data slices from a single theme, despite the latter having three times more training data. This suggests that data diversity has a stronger impact on model performance than data volume alone.
Figure 9: LOSS-based MIAs indicators under different DP.
4.3.6 Comparison of privacy protection intensity
@@
While corpus-level metrics capture direct memorization of training data, they do not fully reflect adversarial privacy risks. To address this limitation, privacy-preserving efficacy is further assessed using the MIA framework duan2024do, which examines whether an adversary can infer training set membership from model outputs. Specifically, samples from $S_{1},S_{2}$} are treated as members, while disjoint samples from $S_{3},S_{4}$} are considered non-members. Throughout federated training, intermediate model checkpoints are evaluated under varying DP budgets $\epsilon\in\{1,3,5,7\}$ and without DP. Both ROC AUC and PR AUC are reported to characterize changes in membership distinguishability. ROC AUC measures the probability that an attacker ranks a randomly chosen member sample higher than a non-member, while PR AUC emphasizes performance under class imbalance, which is common in realistic attack scenarios.
Table 8: Privacy protection effectiveness under different DP budgets across three MIAs. Lower ROC AUC and PR AUC indicate stronger privacy protection. Best results are shown in bold, second-best are underlined.
@@ 5 Conclusion
Despite the effectiveness of the proposed approach, the current privacy protection mechanism is implemented at the client level, and stronger guarantees can be achieved by incorporating supplementary techniques such as active forgetting blanco2025digital and model unlearning jang-etal-2023-knowledge to further mitigate potential privacy risks. Future work will focus on extending MindChat to a multi-modal mental health framework that incorporates speech, facial expressions, and other non-verbal signals to better reflect real-world counseling interactions. Key challenges include effective multi-modal alignment and decoupled representation learning to improve the interpretability and robustness of fusion decisions, as well as fine-grained multi-modal feature extraction and enhancement, combined with data augmentation strategies to strengthen emotional cue modeling.
@@ Appendix A Details of mental support dataset construction
This appendix introduces the construction process of the psychological support dataset in detail. It first introduces the roles and functionalities of each agents involved in the proposed dual-loop multi-agent framework. Then, the prompt templates used to instantiate each agent are presented. Finally, the structured help-seeking process adopted by the seeker agent is elaborated to illustrate how realistic counseling dialogues are generated.
A.1 Description of agents
  • Extractor: The primary function of the extractor is to extract information relevant to the seeker. Raw text data about the seeker collected from public platforms often contains irrelevant interfering information. Moreover, explicit goal setting is instrumental in emotional support conversation Zhou_ACL_23. Consequently, the extractor focuses on extracting key information, encompassing the character, plight, and demand of the seeker, which can be categorized as “who”, “what”, and “how” egarn2012skilled, Hill_helping_20.
  • Seeker: Seeker is tasked with simulating an individual experiencing psychological distress. Its foundational information is derived from the extractor’s output. Given the challenges that LLMs face in mimicking the seeker behaviors Lee_EMNLP_24, we design a multi-stage help-seeking structure to align with the requirements for the seeker in psychological counseling Bohart_APS_2010. Details on this structure can be found in A.3.
  • Supporter: The role played by the supporter is a psychological counselor. The supporter should employ various strategies adapted to different stages of the psychological counseling process Hill_helping_20. Aligned with Liu_ACL_20, we also integrate a combination of stages and strategies to effectively guide the supporter in providing psychological counseling services. However, a distinctive feature of our method is that the strategies adopted by the supporter can be continuously enriched during the interaction between the supporter and the seeker.
A.2 Prompt of agents
The prompt templates used to instantiate the Supporter, Seeker, Evaluator, Corrector, and Manager agents are provided in Figures 10, 11, 12, 13 and 14, respectively.
A.3 Structure of seeking help
The experimental observations demonstrate that the seeker agent often discloses the entire situation in a single statement, contrasting with the gradual self-disclosure process of real-life psychological counseling. To bridge the gap between the seeker agent and the real client, we propose a structure consisting of six stages to facilitate a more natural dialogue between the seeker and the supporter. Table 9 details the six stages.
Table 9: The structure of seeking help.

### v2

@@ IV Experiments
IV-A Baselines
For the dataset-level comparison, we evaluate MindCorpus against seven existing emotional dialogue datasets. These include four Chinese datasets: PsyDTCorpus [40], SoulChatCorpus [8], SMILECHAT [32], and CPsyCounD [47], as well as three English datasets: ExTES [52], ESD-CoT [49], and AUGESC [51].
At the model level, MindChat is benchmarked against a diverse set of LLMs. This set includes widely used general-purpose models such as ChatGPT [1], DeepSeek [26], and Gemini [36], the base model Qwen3-8B [42], and four specialized mental health LLMs SoulChat2.0 [8], CPsyCounX [47], EmoLLM2.0 [43], and MeChat [32].
IV-B Experimental Settings
Figure 4: Data distribution of MindCorpus.
@@ IV-B Experimental Settings
IV-C Results
IV-C1 Synthesized Data Analysis
The synthesized dataset MindCorpus contains 5.75.7k dialogue sessions across diverse themes, generated from curated scenario scripts using a multi-agent collaborative framework. As shown in Fig. 4, dialogues are distributed across multiple themes. Table II summarizes the basic statistics of MindCorpus in comparison with existing emotional dialogue datasets.
TABLE II: Basic Statistics of Emotional Dialogue Datasets.
@@ IV-C1 Synthesized Data Analysis
TABLE III: Automatic and Human Evaluations of Emotional Dialogue Datasets. Highest Scores are in bold.
@@
TABLE IV: Spearman Rank Correlations between Automatic and Human Evaluations across Emotional Dialogue Datasets, Reported per Evaluation Dimension.
@@
TABLE V: Evaluation Results for Data Synthesized by Heterogeneous Multi-agent Groups. Highest Scores are in bold, Second-highest are underlined.
@@
To evaluate data quality, we follow the sampling protocol of [52], randomly selecting $50$ sessions from each dataset while ensuring diversity by incorporating data from clearly classified content categories. We employ GPT-4o as a judge model to score dialogues based on our five-dimensional evaluation framework. Table III reports the results of both automated and human evaluations. The results of automatic evaluation show that MindCorpus attain the best performance on all metrics except for Tru. In human evaluation, aside from Emo., the remaining four metrics consistently indicate that MindCorpus outperforms other emotional dialogue datasets. Overall, both automated and human evaluations confirm the superior quality of MindCorpus compared to existing emotional dialogue datasets. To further examine the consistency between automatic and human evaluations, Spearman rank correlations [46] are computed for each evaluation dimension at the dataset level. As shown in Table IV, most dimensions exhibit moderate to high agreement, demonstrating strong consistency between automated and human evaluations and supporting the reliability of the reported results in Table III.
IV-C2 Heterogeneous Multi-LLM Coordination
We examine the performance of heterogeneous LLMs collaborating within the proposed multi-agent framework, including models of varying sizes and different developers. Table V presents the performance of these LLMs working together in $11$ coordination groups.
During the experiments, we observe that excessively long model outputs tend to exhibit content homogenization, which deviates from real-world psychological counseling scenarios. Empirical statistics of publicly available emotional and counseling dialogue datasets in Table II indicate that the average length of a single response is generally below $100$ words. Motivated by this distribution, we treat dialogues with single-round response lengths under $100$ words as qualified data. Accordingly, Groups $6$ to $11$ in Table V meet this criterion.
With different collaboration groups, the proportion of dialogues requiring modification changes, implying various interactions between agents. More interactions typically result in longer durations. As shown in Groups 5 and 6, compared to Qwen2.5-3B-Instruct, MiniCPM-4B produces fewer qualified dialogues and generates relatively longer responses. However, GPT-4o-based evaluation scores are higher for longer responses, possibly because the automated judges evaluate responses using point-based criteria, whereby longer outputs are more likely to satisfy multiple scoring aspects. In general, larger model sizes correlate with higher data quality, though this trend is not strictly monotonic. In Groups 10 and 11, we employ commercial models for data synthesis via API calls. The synthesized data quality in Group 11 surpasses that of all other experimental groups. However, Group 10 underperforms relative to the small-model collaborations in Groups 8 and 9. This discrepancy may be attributed to performance differences between DeepSeek-V3 and Qwen-Max. As shown in Table V, the quality of data generated through multi-agent collaboration depends not only on model scale but also strongly on model type. Notably, relatively smaller-scale models also demonstrate the potential to achieve performance comparable to that of commercial-grade models.
@@ IV-C2 Heterogeneous Multi-LLM Coordination
IV-C3 Impact of Key Agents on Data Quality
@@
IV-C4 Comparison of Different Psychological LLMs
@@
TABLE VI: Automatic and Human Evaluations of Multiple LLMs in Psychological Counseling Capabilities.
@@
In order to evaluate the performance differences between MindChat and other benchmark models, we conduct both automatic and human evaluations. The results are summarized in Table VI. In the automatic evaluation, MindChat achieves the highest scores across all metrics, better than the general LLMs and the existing mental health-focused LLMs. It is worth noting that MindChat demonstrates superior performance in the dimensions of Comprehensiveness, Professionalism, and Authenticity. These three metrics reflect the depth and quality of psychological support provided, thus highlighting the enhanced capability of MindChat in addressing the psychological background of seeker, offering professional guidance, and establishing authentic empathetic connections. In the human evaluation, DeepSeek and Gemini, as general LLMs with significantly larger parameter sizes, occupy the top two respectively. Despite this disadvantage in terms of model scale, MindChat still ranks third overall. More importantly, it significantly outperforms other models of comparable size, including both general-purpose and domain-specific psychological LLMs, underscoring its effectiveness and competitiveness in mental health scenarios while preserving user privacy.
To further assess the consistency between automatic and human evaluations at the model level, Spearman rank correlations are computed for each evaluation dimension. The results show that Comprehensiveness achieves a high correlation of $0.730$ with a p-value of $0.026$, while Professionalism and Authenticity exhibit lower correlations of $0.351$ with a p-value of $0.354$ and $0.378$ with a p-value of $0.316$, respectively. The observed correlations show clear consistency between automated and human evaluations in Comprehensiveness, whereas Professionalism and Authenticity exhibit relatively weaker alignment. By offering complementary yet distinct perspectives, the two evaluation schemes jointly strengthen the credibility of the model comparison results reported in Table VI.
IV-C5 Effects of Data Quantity and Diversity
In FL architecture, an increasing number of participating clients often leads to larger data quantities and potentially greater diversity. This section investigates how the scale and diversity of training data affect the performance of MindChat under the designed training framework. To evaluate the impact of data quantity, we sample 600 sessions from the Depression and Anxiety dataset and partition them into six equal subsets, as shown in Fig. 7a. For data diversity analysis, we collect 100 samples from each of six distinct themes, outlined in Fig. 7b. All experiments are conducted with $\epsilon=1$, and model performance is compared across settings with $2$, 44, and $6$ collaborating clients. A local baseline is trained using $100$ samples from a single theme.
(a) Quantity slices under same theme.
(b) Theme slices with same quantity.
Figure 7: Comparison of slices: (a) Same theme, (b) Different themes.
TABLE VII: Impact of Data Quantity and Diversity on the Performance of MindChat.
@@ IV-C5 Effects of Data Quantity and Diversity
As shown in Table VII, FL enhances model generalization ability under data isolation, demonstrating the benefits of multi-party collaboration. When training within the same theme, performance increases with the number of clients, indicating the benefit of larger data volume. A similar trend is observed in cross-theme settings, where both data quantity and thematic diversity contribute to performance gains. Notably, models trained with diverse themes outperform those trained with increased data quantity from a single theme. For instance, MindChat-T2, which uses two distinct themes, achieves a higher average score than MindChat-Q6, trained on six data slices from a single theme, despite the latter having three times more training data. This suggests that data diversity has a stronger impact on model performance than data volume alone.
IV-C6 Comparison of Privacy Protection Intensity
@@
(a) ROC AUC under LOSS attack.
(b) PR AUC under LOSS attack.
Figure 9: LOSS-based MIAs indicators under different DP.
TABLE VIII: Privacy Protection Effectiveness under Different DP Budgets across Three MIAs. Lower ROC AUC and PR AUC Indicate Stronger Privacy Protection. Best Results are Shown in bold, Second-best are underlined.
@@
While corpus-level metrics capture direct memorization of training data, they do not fully reflect adversarial privacy risks. To address this limitation, privacy-preserving efficacy is further assessed using the MIA framework [12], which examines whether an adversary can infer training set membership from model outputs. Specifically, samples from $S_{1},S_{2}$} are treated as members, while disjoint samples from $S_{3},S_{4}$} are considered non-members. Throughout federated training, intermediate model checkpoints are evaluated under varying DP budgets $\epsilon\in\{1,3,5,7\}$ and without DP. Both ROC AUC and PR AUC are reported to characterize changes in membership distinguishability. ROC AUC measures the probability that an attacker ranks a randomly chosen member sample higher than a non-member, while PR AUC emphasizes performance under class imbalance, which is common in realistic attack scenarios.
@@ V Conclusion
Despite the effectiveness of the proposed approach, the current privacy protection mechanism is implemented at the client level, and stronger guarantees can be achieved by incorporating supplementary techniques such as active forgetting and model unlearning to further mitigate potential privacy risks. Future work will focus on extending MindChat to a multi-modal mental health framework that incorporates speech, facial expressions, and other non-verbal signals to better reflect real-world counseling interactions. A key direction is to enhance empathetic competence by accurately capturing fine-grained emotional dynamics from multi-modal conversations and strengthening textual representations through data enhancement strategies [25, 29]. Furthermore, improving the discrimination of subtle emotional states via contrastive or adversarial training may enable more effective emotional support, while retrieval-augmented generation can facilitate more personalized and context-aware guidance [21, 2].
@@ Appendix A Details of mental support dataset construction
This appendix introduces the construction process of the psychological support dataset in detail. It first introduces the roles and functionalities of each agents involved in the proposed dual-loop multi-agent framework. Then, the structured help-seeking process adopted by the seeker agent is elaborated to illustrate how realistic counseling dialogues are generated.
TABLE IX: The Structure of Seeking Help.
@@
A-A Description of agents
  • Extractor: The primary function of the extractor is to extract information relevant to the seeker. Raw text data about the seeker collected from public platforms often contains irrelevant interfering information. Moreover, explicit goal setting is instrumental in emotional support conversation [53]. Consequently, the extractor focuses on extracting key information, encompassing the character, plight, and demand of the seeker, which can be categorized as “who”, “what”, and “how” [13, 17].
  • Seeker: Seeker is tasked with simulating an individual experiencing psychological distress. Its foundational information is derived from the extractor’s output. Given the challenges that LLMs face in mimicking the seeker behaviors [23], we design a multi-stage help-seeking structure to align with the requirements for the seeker in psychological counseling [3]. Details on this structure can be found in Appendix A-C.
  • Supporter: The role played by the supporter is a psychological counselor. The supporter should employ various strategies adapted to different stages of the psychological counseling process [17]. Aligned with [27], we also integrate a combination of stages and strategies to effectively guide the supporter in providing psychological counseling services. However, a distinctive feature of our method is that the strategies adopted by the supporter can be continuously enriched during the interaction between the supporter and the seeker.
A-B Prompt of agents
The prompt templates used to instantiate the Supporter, Seeker, Evaluator, Corrector, and Manager agents are provided in Figures 10, 11, 13, 14 and 12, respectively.
A-C Structure of seeking help
The experimental observations demonstrate that the seeker agent often discloses the entire situation in a single statement, contrasting with the gradual self-disclosure process of real-life psychological counseling. To bridge the gap between the seeker agent and the real client, we propose a structure consisting of six stages to facilitate a more natural dialogue between the seeker and the supporter. Table IX details the six stages.